# Causal and Stable Superfluid Hydrodynamics

### v1

@@ 1 Introduction
This paper is organized as follows. In section 2, we give a detailed overview of first order relativistic superfluid hydrodynamics. We elucidate the state of thermal equilibrium, present the general first order constitutive relations when out of equilibrium, discuss frame transformations and frame invariants, and finally introduce the LLCP frame. Then, in section 3, we take a closer look at the dynamics in the LLCP frame and argue that it is unstable and acausal. Next, in section 4, working in a general hydrodynamic frame, we enunciate the criteria transport parameters must satisfy to ensure linear stability of the equilibrium state (subsection 4.1), followed by laying out the criteria for linear causality of propagation for perturbations about the equilibrium state (subsection 4.2). We conclude with a discussion and outlook in section 5. Appendix A contains additional details about the LLCP frame and the physical transport coefficients, while appendix B summarizes the mathematical criteria that need to be imposed for ensuring stability and causality of the equilibrium state.
@@ 2.1 Thermal equilibrium
When the global $U(1)$ symmetry gets spontaneously broken, the system enters a superfluid phase. Microscopically, the spontaneous breakdown happens because an operator charged under the symmetry condenses and acquires a non-vanishing expectation value in the equilibrium state. The phase of the condensed operator, denoted $\varphi$, emerges as an additional degree of freedom for the system - it is the Goldstone mode associated with the spontaneous breaking of the global $U(1)$ symmetry. Under the gauge transformation $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\lambda$, it undergoes a shift via $\varphi\rightarrow\varphi+\lambda$. In thermal equilibrium, we demand that the Goldstone $\varphi$ is invariant under a combination of diffeomorphisms generated by $K^{\mu}$ and gauge transformations generated by $\Lambda_{K}$, via
@@
with $F_{\mu\nu}\equiv\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ being the field strength tensor for the gauge field $A_{\mu}$. In the hydrodynamic regime, while the derivatives of $\varphi$ are not necessarily small, the derivatives of $\xi^{\mu}$ are small, justifying its use as the appropriate degree of freedom Bhattacharya:2011eea . Notice that $\xi^{\mu}$ has a well-defined microscopic definition, just as the energy-momentum tensor and the $U(1)$ current. In what follows, we will treat $\xi^{\mu}$ as an independent (non-conserved) current, [^4] This stands in contrast with the usual treatment of $\xi^{\mu}$ in the literature Herzog:2011ec ; Bhattacharya:2011eea ; Bhattacharya:2011tra ; Jensen:2012jh ; Bhattacharyya:2012xi , where it is treated as a hydrodynamical variable characterizing the equilibrium state at par with $T,\mu$ etc. Essentially, in our approach, we switch the side of the derivative expansion on which $\xi^{\mu}$ sits. at par with $T^{\mu\nu}$ and $J^{\mu}$, whose evolution is constrained by eq. (4).
@@ 2.3 Frame transformations and frame invariants
In the conformal limit, the tracelessness of the energy-momentum tensor leads to the relation $\mathcal{E}_{1}-\mathcal{V}_{1}=(d-1)\mathcal{P}_{1}$, implying $\mathcal{F}=-(d-1)\mathcal{G}$, and thus leads to one less frame invariant scalar at $\mathcal{O}(\partial)$.
@@ 3.3 Acausality and instability of the LLCP frame
It turns out that the LLCP frame also violates the other two conditions required for causality. In the limit $|\mathbf{k}|\to\infty$, there are are only five modes that go linearly in $|\mathbf{k}|$. Of the remaining four modes, one behaves at $|\mathbf{k}|\to\infty$ as $\omega\sim|\mathbf{k}|^{3}$, while the remaining three take the form $\omega\sim|\mathbf{k}|^{4/3}$. These violate the first two conditions of eq. (55). The LLCP frame is thus acausal, and therefore there also exists a boosted reference frame in which it is linearly unstable Gavassino:2021owo ; Gavassino:2023myj . These observations motivate one to look for hydrodynamic frames where the out of equilibrium dynamics of superfluids remains stable and causal, a subject which we turn next to.
@@ 5 Discussion and outlook
Several straightforward extensions of the ideas discussed in this paper immediately present themselves. Firstly, the superfluids we have discussed are on a flat background; it would be very interesting to extend the discussion to superfluids coupled to dynamical gravity. The cores of neutron stars are posited to be superfluidic in nature MIGDAL1959655 ; BAYM1969 , and therefore a general relativistic causal theory seems relevant. Secondly, a more detailed description of superfluids should also include the effects of vortices.
Finally, in this paper we performed a linearized analysis of the equations of motion for a relativistic superfluid. An important outstanding question is the classification of the non-linear system of partial differential equations. Are the BDNK superfluid equations strongly hyperbolic? Do they admit a well-posed initial value problem? What is the role of the constraint equations present in (35c)? Following the discussion in Hoult:2024mhd , it is likely that a linearized analysis is sufficient to determine non-linear causality of the equations of motion; however, in the present paper we explicitly solved the constraint equations. The effect this has on the equivalence between linear and non-linear causality described in Hoult:2024mhd remains to be seen. The causality of relativistic magnetohydrodynamics Armas:2022wvb ; Hoult:2024mhd is likely a useful guide, as it is also a system containing constraint equations (namely, $\nabla{\cdot}B=0$).
@@ A.1 Positivity of entropy production
We next impose the Onsager reciprocity condition,191919 The Onsager relations are the condition that retarded two-point Green’s functions are covariant under time-reversal, see e.g. Kovtun:2012rj . which leads to the demand that the coefficients of the cross terms between independent scalars and vectors in the expression above be identical. This immediately leads to the relations given in eq. (48) of the main text, leading to a reduction in the number of independent transport parameters from 14 to 10. The remaining 10 parameters form the set of physical transport coefficients for the conformal superfluid at $\mathcal{O}(\partial)$ in the derivative expansion. These are further constrained by the requirement of positive semidefiniteness of entropy production i.e., $\partial_{\mu}J^{\mu}_{S}\geq 0$. Mathematically, this translates to ensuring that the RHS of eq. (89) is a positive semidefinite quadratic form. This in turn leads to the constraints quoted in eq. (49) on the LLCP frame transport coefficients.

### v3

@@ 1 Introduction
This paper is organized as follows. In section 2, we give a detailed overview of first order relativistic superfluid hydrodynamics. We elucidate the state of thermal equilibrium, present the general first order constitutive relations when out of equilibrium, discuss frame transformations and frame invariants, and finally introduce the LLCP frame. Then, in section 3, we take a closer look at the dynamics in the LLCP frame and argue that it is unstable and acausal. Next, in section 4, working in a general hydrodynamic frame, we enunciate the criteria transport parameters must satisfy to ensure linear stability of the equilibrium state (subsection 4.1), followed by laying out the criteria for linear causality of propagation for perturbations about the equilibrium state (subsection 4.2). We conclude with a discussion and outlook in section 5. Appendix A contains additional details about the LLCP frame and the physical transport coefficients, while appendix B summarizes the mathematical criteria that need to be imposed for ensuring stability and causality of the equilibrium state. In appendix C, we examine the “superfluid Landau frame” of Arean:2023nnn , which we dub the modified Landau frame, and argue that the dynamics in this frame too is acausal and unstable, as in the LLCP frame.
@@ 2.1 Thermal equilibrium
When the global $U(1)$ symmetry gets spontaneously broken, the system enters a superfluid phase. Microscopically, the spontaneous breakdown happens because an operator charged under the symmetry condenses and acquires a non-vanishing expectation value in the equilibrium state. The phase of the condensed operator, denoted $\varphi$, emerges as an additional degree of freedom for the system - it is the Goldstone mode associated with the spontaneous breaking of the global $U(1)$ symmetry. [^4] Throughout the paper, we work in the regime far removed from the superfluid critical point, decoupling the gapped “amplitude mode” Pekker-Varma ; Donos:2022xfd from the rest of the dynamics. Under the gauge transformation $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\lambda$, it undergoes a shift via $\varphi\rightarrow\varphi+\lambda$. In thermal equilibrium, we demand that the Goldstone $\varphi$ is invariant under a combination of diffeomorphisms generated by $K^{\mu}$ and gauge transformations generated by $\Lambda_{K}$, via
@@
with $F_{\mu\nu}\equiv\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ being the field strength tensor for the gauge field $A_{\mu}$. It is important to mention here that with regards to derivative counting, one takes the Goldstone mode $\varphi$ to be $\mathcal{O}(\partial^{-1})$, implying that its derivative $\xi^{\mu}$ is an $\mathcal{O}(1)$ quantity. Notice that $\xi^{\mu}$ has a well-defined microscopic definition, just as the energy-momentum tensor and the $U(1)$ current. In what follows, we will treat $\xi^{\mu}$ as an independent (non-conserved) current, [^5] This stands in contrast with an alternate treatment of $\xi^{\mu}$ in parts of the literature e.g. Herzog:2011ec ; Bhattacharya:2011eea ; Bhattacharya:2011tra ; Jensen:2012jh ; Bhattacharyya:2012xi , where it is treated as a hydrodynamical variable characterizing the equilibrium state at par with $T,\mu$ etc. at par with $T^{\mu\nu}$ and $J^{\mu}$, whose evolution is constrained by eq. (4).
@@ 2.3 Frame transformations and frame invariants
In the conformal limit, the tracelessness of the energy-momentum tensor leads to the relation $\mathcal{E}_{1}-\mathcal{V}_{1}=(d-1)\mathcal{P}_{1}$, implying that $\mathcal{F}=-(d-1)\mathcal{G}$,101010 This may be seen straightforwardly by considering the definitions of ${\cal F}$ and ${\cal G}$ in a frame with $\mathcal{E}_{1}=\mathcal{N}_{1}=\mathcal{S}_{1}=\mathcal{Z}_{1}=\mathcal{J}_{1}^{\mu}=\mathcal{X}_{1}^{\mu}=0\,,$ (44) which we will presently introduce as the “LLCP frame” in section 2.4. In this frame, ${\cal F}={\cal V}_{1}^{\rm LLCP}$, ${\cal G}={\cal P}_{1}^{\rm LLCP}$, and the relationship $\mathcal{E}_{1}-\mathcal{V}_{1}=(d-1)\mathcal{P}_{1}$ becomes $-{\cal V}_{1}^{\rm LLCP}=\left(d-1\right){\cal P}_{1}^{\rm LLCP}$. This implies that ${\cal F}={\cal V}_{1}^{\rm LLCP}=-\left(d-1\right){\cal P}_{1}^{\rm LLCP}=-\left(d-1\right){\cal G}$, and hence ${\cal F}=-\left(d-1\right){\cal G}$. However, ${\cal F}$ and ${\cal G}$ are both frame invariant quantities, thereby implying that this relation must be true in any hydrodynamic frame. and thus leads to one less frame invariant scalar at $\mathcal{O}(\partial)$.
@@ 3.3 Acausality and instability of the LLCP frame
It turns out that the LLCP frame also violates the other two conditions required for causality. In the limit $|\mathbf{k}|\to\infty$, there are are only five modes that go linearly in $|\mathbf{k}|$. Of the remaining four modes, one behaves at $|\mathbf{k}|\to\infty$ as $\omega\sim|\mathbf{k}|^{3}$, while the remaining three take the form $\omega\sim|\mathbf{k}|^{4/3}$. These violate the first two conditions of eq. (56). The LLCP frame is thus acausal, and therefore there also exists a boosted reference frame in which it is linearly unstable Gavassino:2021owo ; Gavassino:2023myj . [^19] In appendix C, we take a close look at another important frame, the “superfluid Landau frame” of Arean:2023nnn , which we call the modified Landau frame, and argue that it too is acausal and unstable. These observations motivate one to look for hydrodynamic frames where the out of equilibrium dynamics of superfluids remains stable and causal, a subject which we turn next to.
@@ 5 Discussion and outlook
Further, the superfluids we have discussed in the present work are on a flat background; it would be very interesting to extend the discussion to superfluids coupled to dynamical gravity. The cores of neutron stars are posited to be superfluidic in nature MIGDAL1959655 ; BAYM1969 , and therefore a general relativistic causal theory seems relevant. Secondly, a more detailed description of superfluids should also include the effects of vortices.
Additionally, we have assumed that the superfluid under investigation is in the regime far removed from the critical point, and the dynamics of the “amplitude mode” Pekker-Varma ; Donos:2022xfd can be neglected. It would be interesting to see how the inclusion of the amplitude mode into our analysis, relevant when the system is near criticality, will affect the causality and stability aspects.
Finally, in this paper we performed a linearized analysis of the equations of motion for a relativistic superfluid. An important outstanding question is the classification of the non-linear system of partial differential equations. Are the BDNK superfluid equations strongly hyperbolic? Do they admit a well-posed initial value problem? What is the role of the constraint equations present in (35c)? Following the discussion in Hoult:2024qph , it is likely that a linearized analysis is sufficient to determine non-linear causality of the equations of motion; however, in the present paper we explicitly solved the constraint equations. The effect this has on the equivalence between linear and non-linear causality described in Hoult:2024qph remains to be seen. The causality of relativistic magnetohydrodynamics Armas:2022wvb ; Hoult:2024qph is likely a useful guide, as it is also a system containing constraint equations (namely, $\nabla{\cdot}B=0$).
@@ A.1 Positivity of entropy production
We next impose the Onsager reciprocity condition,222222 The Onsager relations are the condition that retarded two-point Green’s functions are covariant under time-reversal, see e.g. Kovtun:2012rj . which leads to the demand that the coefficients of the cross terms between independent scalars and vectors in the expression above be identical. This immediately leads to the relations given in eq. (49) of the main text, leading to a reduction in the number of independent transport parameters from 14 to 10. The remaining 10 parameters form the set of physical transport coefficients for the conformal superfluid at $\mathcal{O}(\partial)$ in the derivative expansion. These are further constrained by the requirement of positive semidefiniteness of entropy production i.e., $\partial_{\mu}J^{\mu}_{S}\geq 0$. Mathematically, this translates to ensuring that the RHS of eq. (90) is a positive semidefinite quadratic form. This in turn leads to the constraints quoted in eq. (50) on the LLCP frame transport coefficients.
@@ B.2 Schur-Cohn Stability
Appendix C Acausality of the Modified Landau Frame
The LLCP frame discussed in section 2.4 was shown to be acausal, and therefore unstable (see section 3.3). However, there are other important superfluid hydrodynamic frames as well. An example is the so-called “Landau frame” of Arean:2023nnn . To avoid confusion with the LLCP frame, as well as the usual Landau frame used in relativistic hydrodynamics, we will refer to this as the modified Landau frame. This frame is defined by the choices
${\cal E}_{1}={\cal U}_{1}={\cal N}_{1}={\cal Z}_{1}=0,\quad{\cal Q}_{1}^{\mu}={\cal X}_{1}^{\mu}=0\,.$ (103)
This is equivalent to
$u_{\mu}T_{1}^{\mu\nu}=0,\quad u_{\mu}J_{1}^{\mu}=0,\quad\xi_{\mu}=-\left(\mu+{\cal M}_{1}\right)u_{\mu}+\zeta_{\mu}\,.$ (104)
The first-order corrections to the stress-energy tensor, charge current, and phase field are then given by
$T^{\mu\nu}_{1}$ $=-\left(d-1\right){\cal P}_{1}z^{\mu}z^{\nu}+{\cal P}_{1}\tilde{\Delta}^{\mu\nu}+2{\cal R}_{1}^{(\mu}z^{\nu)}+{\cal T}_{1}^{\mu\nu}\,,$ (105a)
$J^{\mu}_{1}$ $={\cal S}_{1}z^{\mu}+{\cal J}^{\mu}_{1}\,,$ (105b)
$\xi^{\mu}$ $=-{\cal M}_{1}u^{\mu}\,.$ (105c)
After demanding positivity of entropy production and imposing the Onsager relations, one finds that the non-vanishing dissipative corrections are given by
${\cal P}_{1}$ $=-\frac{\tilde{\lambda}_{6}T}{d}\mathbb{s}_{4}+\tilde{\pi}_{6}\mathbb{s}_{6}-\frac{T^{d}\tilde{\alpha}_{6}}{d}\mathbb{s}_{7}\,,$ (106a)
${\cal R}_{1}^{\mu}$ $=T\tilde{\gamma}_{3}\mathbbmss{v}_{2}^{\mu}+\tilde{\varrho}_{3}\mathbbmss{v}_{3}^{\mu}\,,$ (106b)
${\cal T}_{1}^{\mu\nu}$ $=-\tilde{\eta}\sigma^{\mu\nu}\,,$ (106c)
${\cal S}_{1}$ $=\tilde{\lambda}_{4}\mathbb{s}_{4}+\tilde{\lambda}_{6}\mathbb{s}_{6}+T^{d-1}\alpha_{4}\mathbb{s}_{7}\,,$ (106d)
${\cal J}^{\mu}_{1}$ $=\tilde{\gamma}_{2}\mathbbmss{v}_{2}^{\mu}+\tilde{\gamma}_{3}\mathbbmss{v}_{3}^{\mu}\,,$ (106e)
${\cal M}_{1}$ $=\tilde{\alpha}_{4}\mathbb{s}_{4}+\tilde{\alpha}_{6}\mathbb{s}_{6}+\tilde{\alpha}_{7}\mathbb{s}_{7}.$ (106f)
Performing a linearized analysis in this frame, which falls into the class of “constraint frames,” eq. (54), we find that there are six hydrodynamic modes, and zero non-hydrodynamic modes, which clearly violates the third condition of eq. (56). Indeed, we find that of the six hydrodynamics modes, five behave as $\omega\sim k^{2}$ in the limit of large $k$, violating the first two conditions of eq. (56). The sixth goes as $\omega\sim z{\cdot}k$. This mode is therefore acausal, and there exists a boosted reference frame in which it will exhibit instability. This instability, unlike the physical instability described in Arean:2023nnn , is observer dependent, and not physical.